OpenStudy (anonymous):

∫∫ (x+y)^2/(X^2+Y^2) dxdy The region bounded by the positive x and y axes and the line y = 1-x

OpenStudy (zarkon):

expand \((x+y)^2\) then simplify...then integrate

OpenStudy (anonymous):

i know how to intefrate this i just dont know how to convert the limits into polar form?

OpenStudy (anonymous):

i know that im left with sin2theta but the integrals of x: 0 to 1 and 0 to 1-y, how can i convert the limits into polar form?

OpenStudy (anonymous):

Can anyone confirm that im on the right lines ?

OpenStudy (zarkon):

\[y = 1-x\] \[r\sin(\theta)=1-r\cos(\theta)\] \[r\sin(\theta)+r\cos(\theta)=1\] \[r(\sin(\theta)+\cos(\theta))=1\] \[r=\frac{1}{\sin(\theta)+\cos(\theta)}\]

OpenStudy (anonymous):

Are the limits r=1, r=0 and theta=pi/4, to theta=0 ?

OpenStudy (anonymous):

so what are the limits ? are mine correct?

OpenStudy (zarkon):

those are not correct

OpenStudy (anonymous):

i sketched a graph and theses are the one that i see.. what im i doing wrong?

OpenStudy (zarkon):



OpenStudy (zarkon):

why do you have r=1 and \(\theta=\pi/4\)

OpenStudy (zarkon):

that makes 1/8th of a circle

OpenStudy (anonymous):

polar coordinates have to be with respect to rdrdtheta

OpenStudy (zarkon):

yes...I'm questioning the choice of 1 and pi/4

OpenStudy (zarkon):

you are integrating over a triangle in the 1st quadrant

OpenStudy (zarkon):

not a circle

OpenStudy (anonymous):

i think r =1 to r=0 is correct but not sure about theta

OpenStudy (zarkon):

it is not

OpenStudy (anonymous):

hmmm let me see...

OpenStudy (zarkon):